OpenStudy (anonymous):

Of the following options, what could be a possible first step in solving the equation 3x − 1 = 2x + 4? Combining like terms, 3x + 2x = 5x Subtracting 1 from both sides of the equation Subtracting 2x from both sides of the equation Adding 4 to both sides of the equation

OpenStudy (anonymous):

@Olyalhic

OpenStudy (anonymous):

C ?

OpenStudy (texaschic101):

yes..it is C

OpenStudy (olyalhic):

@gabylovesyou100 yes its c

OpenStudy (texaschic101):

great minds think alike :)

OpenStudy (anonymous):

thank you

OpenStudy (olyalhic):

@gabylovesyou100 heres the steps for it just so you know subtract 2x from both sides 3x−1−2x=2x+4−2x x−1=4 Add 1 to both sides. x−1+1=4+1 x=5 Answer x=5

OpenStudy (anonymous):

:) ty !